OpenStudy (anonymous):

Anyone mind helpin me with this? Determine which system below will produce infinitely many solutions. -4x + 6y = 20 4x - 3y = -2 -2x + 4y = 18 4x - 8y = -36 3x - 2y = 5 -6x - 4y = -10 5x - 3y = 12 15x - 9y = 24

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

The 2nd one has infinite solutions

OpenStudy (anonymous):

how can you tell?

OpenStudy (anonymous):

whats the work you did to figure that out i mean ?

OpenStudy (anonymous):

nvm i figured it out, thanks love

OpenStudy (anonymous):

Use the system of equations to get an absolute truth. Example: y = x+5 2y = 2x+10 If you solve this you get: x+5 = x+5 5 = 5 If you get to a statement that is obviously false, then you have no solutions. For example: y=9x+5 y=9x-2 Solving these you get to 5=-2 So there are no solutions to that. Do you understand?

OpenStudy (khally92):